If frank is 7 years older than chuck and the sum of their ages was 15 how old is chuck?
Answer by algebrahouse.com(1659)

x = Chuck's age
x + 7 = Frank's age

x + x + 7 = 15 {the sum of their ages is 15}
2x + 7 = 15 {combined like terms}
2x = 8 {subtracted 7 from each side}
x = 4 {divided each side by 2}

Chuck is 4

For more help from me, visit: www.algebrahouse.com